OpenStudy (anonymous):

Use a ratio to compare the values. Simplify the ratio. 12 yd to 20 ft A. 5 : 9 B. 3 : 5 C. 5 : 3 D. 9 : 5

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

@friendzone

jhonyy9 (jhonyy9):

sorry to the first one is D. the right answer because 24 to 30 what simplified add 4 to 5

jhonyy9 (jhonyy9):

ok.?

OpenStudy (anonymous):

wait its 4 to 5

jhonyy9 (jhonyy9):

so to this 12/20 = 3*4/4*5 simplifie it by 4 and what will remaine there ?

jhonyy9 (jhonyy9):

sorry

OpenStudy (anonymous):

15 to 11

OpenStudy (anonymous):

i got 11

jhonyy9 (jhonyy9):

there are yard and feet so first you need to change them or in yard or in feet - ok. ?

OpenStudy (anonymous):

when i did 12/20 = 3*4/4*5 simplifie it by 4 and what will remaine there

OpenStudy (anonymous):

i got 11

OpenStudy (anonymous):

so how do you turn yards into feet

OpenStudy (anonymous):

?

jhonyy9 (jhonyy9):

you dont know it ?

jhonyy9 (jhonyy9):

1 yard = 3 feet

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so than 12 yard = 36 feet this mean 36 feet to 20 ft - make the ratio

OpenStudy (anonymous):

oh now i get it kinda

jhonyy9 (jhonyy9):

36/20 = 6*5/5*4 = 6/4 = 3/2 yes ?

jhonyy9 (jhonyy9):

sorry my mistaque 36/20 9/5

OpenStudy (anonymous):

is it A

OpenStudy (anonymous):

oh so D?

jhonyy9 (jhonyy9):

not A. is D. sure

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

ok so D

jhonyy9 (jhonyy9):

was my pleasure - good luck

OpenStudy (anonymous):

ty

jhonyy9 (jhonyy9):

you are welcome anytime

OpenStudy (anonymous):

can you help me with 2 more

OpenStudy (anonymous):

or no

jhonyy9 (jhonyy9):

yes but write them on the open question column please

OpenStudy (anonymous):

what is a question column

jhonyy9 (jhonyy9):

open questions

jhonyy9 (jhonyy9):

where you have wrote this

OpenStudy (anonymous):

oh do i make a new one

jhonyy9 (jhonyy9):

and tag me there - ok. ?